# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v7

@@ III-B Real-Scene Based Simulators
Achieving universal embodied agents in household activities has been a primary focus in the field of embodied AI research. These embodied agents need to deeply understand human daily life and perform complex embodied tasks such as navigation and interaction in indoor environments. To meet the demands of these complex tasks, the simulated environments need to be as close to real world as possible, which places high demands on the complexity and realism of the simulators. This led to the creation of simulators based on real world environments. These simulators mostly collect data from the real world, create photorealistic 3D assets, and build scenes using 3D game engines like UE5 and Unity. The rich and realistic scenes make simulators based on real world environments the top choice for research on embodied AI in household activities.
@@
AI2-THOR [61] is an indoor embodied scene simulator based on Unity3D, led by the Allen Institute for Artificial Intelligence. As a high-fidelity simulator built in the real world, AI2-THOR has richly interactive scene objects and the physical properties assigned to them (such as open/close or even cold/hot). AI2-THOR consists of two parts: iTHOR and RoboTHOR. iTHOR contains 120 rooms categorized as kitchens, bedrooms, bathrooms, and living rooms, with over 2000 unique interactive objects, and supports multi-agent simulation; RoboTHOR contains 89 modular apartments with 600+ objects, the uniqueness of which is that these apartments correspond to real scenes in the real world. So far, more than a hundred works have been published based on AI2-THOR.
Matterport 3D [62] is proposed in R2R[63] and is more commonly used as a large-scale 2D-3D visual dataset. The Matterport3D dataset includes 90 architectural indoor scenes, comprises 10,800 panoramas and 194,400 RGB-D images, and provides surface reconstruction, camera posture, and 2D and 3D semantic segmentation annotations. Matterport3D transforms 3D scenes into discrete “viewpoints”, and embodied agents move between adjacent “viewpoints” in Matterport3D scenes. At each “viewpoint”, embodied agents can obtain a 1280x1024 panorama image (18$\times$ RGB-D) centered on the “viewpoint”. Matterport3D is one of the most important embodied navigation benchmarks.
Virtualhome [64] is a home activity embodied AI simulator brought by Puig et al. What makes Virtualhome special most is its environment represented by an environment graph. The environment graph represents the objects in the scene and their related relationships. Users can also customize and modify the environment graph to achieve the custom configuration of scene objects. This kind of environment graph provides a new way for embodied agents to understand the environment. Similar to AI2-THOR, Virtualhome also provides a large number of interactive objects, and embodied agents can interact with them and change their status. Another feature of Virtualhome is its simple and easy-to-use API. The actions of embodied agents are simplified to the format of “operation + object”. This feature makes Virtualhome widely used in the research fields of embodied planning, instruction decomposition, etc.
Habitat [65] is an open-source simulator for large-scale human-robot interaction launched by Meta. Based on the Bullet physics engine, Habitat has implemented high-performance, high-speed, parallel 3D simulation, and provides a rich interface for reinforcement learning of embodied agents. Habitat has extremely high degree of openness. Researchers can import and create 3D scenes in Habitat or use the rich open resources on the Habitat platform for expansion. Habitat has many customizable sensors and supports multi-agent simulation. Multiple embodied agents from open resources or customizations (e.g., humans and robot dogs) can cooperate in the simulator, move freely, and perform simple interactions with the scene. Thus, Habitat is attracting increasing attention.
TABLE III: Comparison of Real-Scene Based Simulators. For the sensor, S refers to Semantic, L refers to LiDAR and A refers to Audio.
Simulation Platform Year Num of Scenes Continuous Action 3D Scene Scans Sensors Physics Multiple Agents Object States 3D Assets
[…]
TDW [69] 2021 - $\surd$ $\surd$ RGB-D, S, A $\surd$ $\surd$ $\surd$ $\surd$
Different from other simulators that focus more on the scene, SAPIEN [66] pays more attention to simulate the interaction of objects. Based on the PhysX physics engine, SAPIEN provides fine-grained embodied control, which can implement joint control based on force and torque through the ROS interface. Based on the PartNet-Mobility Dataset, SAPIEN provides indoor simulation scenes containing rich interactive objects and supports the import of custom resources. Different from simulators like AI2-THOR that directly changes the status of objects, SAPIEN supports simulated physical interactions, and embodied agents can control the hinged parts of objects through physical actions, thereby changing the status of objects. These features make SAPIEN very suitable for training the fine-grained object operation of embodied AI.
iGibson [67][68] is an open-source simulator launched by Stanford. Built on the Bullet physics engine, iGibson provides 15 high-quality indoor scenes, and supports the import of assets from other datasets such as Gibson and Matterport3D. As an object-oriented simulator, iGibson assigns rich changeable attributes to objects, not limited to the kinematic properties of objects (posture, speed, acceleration, etc.), but also includes temperature, humidity, cleanliness, switch status, etc. In addition, besides the depth and semantic sensors that are standard in other simulators, iGibson also provides LiDAR for embodied agents, allowing agents to obtain 3D point clouds in the scene easily. Regarding embodied agent configuration, iGibson supports continuous action control and fine-grained joint control. This allows the embodied agents in iGibson to interact delicately with the objects while moving freely.
TDW [69] was launched by MIT. As one of the latest embodied simulators, TDW combines high-fidelity video and audio rendering, realistic physical effects, and a single flexible controller, making certain progress in the perception and interaction of the simulated environment. TDW integrates multiple physics engines into one framework, which can realize the physical interaction simulation of various materials such as rigid bodies, soft bodies, fabrics, and fluids, and provides situational sounds when interacting with objects. Thus, TDW has taken an important step compared to other simulators. TDW supports the deployment of multiple intelligent agents and provides users with a rich API library and asset library, allowing users to freely customize scenes and tasks according to their own needs, even outdoor scenes and related tasks.
@@
Table III summarizes all the simulators based on the real scenarios mentioned above. Sapien stands out for its design, specifically tailored to simulate interactions with joint objects like doors, cabinets, and drawers. VirtualHome is notable for its unique environment graph, which facilitates high-level embodied planning based on natural language descriptions of environments. While AI2Thor offers a wealth of interactive scenes, these interactions, similar to those in VirtualHome, are script-based and lack real physical interactions. This design suffices for embodied tasks not requiring fine-grained interactions. Both iGibson and TDW provide fine-grained embodied control and highly simulated physical interactions. iGibson excels in offering abundant and realistic large-scale scenes, making it suitable for complex and long-term mobile operations, whereas TDW allows greater user freedom in scene expansion and features unique audio and flexible fluid simulations, making it indispensable for related simulation scenarios. Matterport3D, a foundational 2D-3D visual dataset, is widely used and extended in embodied AI benchmarks. Although the embodied agent in Habitat lacks interaction capabilities, its extensive indoor scenes, user-friendly interfaces, and open framework make it highly regarded in embodied navigation.
@@
Besides, automated simulation scene construction is greatly beneficial for obtaining high-quality embodied data. RoboGen [70] customizes tasks from randomly sampled 3D assets through LLMs, thereby creating scenes and automatically training agents; HOLODECK [71] can automatically customize corresponding high-quality simulation scenes in AI2-THOR based on human instructions; PhyScene [72] generates interactive and physically consistent high-quality 3D scenes based on conditional diffusion. The Allen Institute for Artificial Intelligence expanded AI2-THOR and proposed ProcTHOR [73], which can automatically generate simulated scenes with sufficient interactivity, diversity, and rationality. These methods provide insights into the embodied AI.
@@ IV-A Active Visual Perception
TABLE IV: The comparison of the active visual perception methods.
@@
vSLAM Traditional vSLAM MonoSLAM[75], MSCKF[76], PTAM[77],
ORB-SLAM[78], DTAM[79], LSD-SLAM[80]
Semantic vSLAM SLAM++[81], CubeSLAM[82], HDP-SLAM[83], QuadricSLAM[84], So-SLAM[85],
DS-SLAM[86], DynaSLAM[87], SG-SLAM[88], OVD-SLAM[89], GS-SLAM[90]
@@
PT3[105], 3D-VisTA[106], LEO[107], PQ3D[108], PointMamba[109], PCM[110], Mamba3D[111]
@@
Active visual perception systems require fundamental capabilities such as state estimation, scene perception, and environment exploration. As shown in Fig. 7, these capabilities have been extensively studied within the domains of Visual Simultaneous Localization and Mapping (vSLAM)[118, 119], 3D Scene Understanding[120], and Active Exploration[12]. These research areas contribute to developing robust active visual perception systems, facilitating improved environmental interaction and navigation in complex, dynamic settings. We briefly introduce these three components and summarize the methods mentioned in each part in Table IV.
@@
Figure 7: The schematic diagram of active visual perception. Visual SLAM and 3D Scene Understanding provide the foundation for passive visual perception, while active exploration can provide activeness to the passive perception system. These three elements complement each other and are essential to the active visual perception system.
IV-A1 Visual Simultaneous Localisation and Mapping
Simultaneous Localization and Mapping (SLAM) is a technique that determines a mobile robot’s position in an unknown environment while concurrently constructing a map of that environment[121, 122]. Range-based SLAM [123, 124, 125] creates point cloud representations using rangefinders (e.g., laser scanners, radar, and/or sonar), but is costly and provides limited environmental information. Visual SLAM (vSLAM)[118, 119] uses on-board cameras to capture frames and construct a representation of the environment. It has gained popularity due to its low hardware cost, high accuracy in small-scale scenarios, and ability to capture rich environmental information. Classical vSLAM techniques can be divided into Traditional vSLAM and Semantic vSLAM [119].
Traditional vSLAM systems estimate the robot’s pose in an unknown environment using image information and multi-view geometry principles to construct a low-level map (e.g., sparse maps, semi-dense maps, and dense maps) composed of point clouds, such as filter-based methods (e.g., MonoSLAM [75], MSCKF [76]), keyframe-based methods (e.g., PTAM [77], ORB-SLAM [78]), and direct tracking methods (e.g., DTAM [79], LSD-SLAM [80]). Since point clouds in low-level maps do not correspond directly to objects in the environment, making them difficult for embodied robots to interpret and utilize. However, the advent of semantic concepts, particularly semantic vSLAM systems integrated with semantic information solutions, has significantly improved robots’ ability to perceive and navigate unexplored environments.
Early works, such as SLAM++[81], use real-time 3D object recognition and tracking to create efficient object graphs, enabling robust loop closure, relocalization, and object detection in cluttered environments. CubeSLAM[82] and HDP-SLAM[83] introduce 3-D rectangular into the map to construct a lightweight semantic map. QuadricSLAM[84] employs semantic 3D ellipsoids to achieve precise modeling of object shapes and poses in complex geometrical environments. So-SLAM[85] incorporates fully coupled spatial structure constraints (coplanarity, collinearity, and proximity) in indoor environments. To meet the challenges of dynamic environments, DS-SLAM[86], DynaSLAM[87] and SG-SLAM[88] employ semantic segmentation for motion consistency checks and multiview geometry algorithms to identify and filter dynamic objects, ensuring stable localization and mapping. OVD-SLAM[89] leverages semantic, depth, and optical flow information to distinguish dynamic regions without predefined labels, achieving more accurate and robust localization. GS-SLAM[90] utilizes 3D Gaussian representation that balances efficiency and accuracy through a real-time differentiable splatting rendering pipeline and adaptive expansion strategy.
@@ IV-A1 Visual Simultaneous Localisation and Mapping
3D scene understanding aims to distinguish objects’ semantics, identify their locations, and infer the geometric attributes from 3D scene data, which is fundamental in autonomous driving[126], robot navigation[127], and human-computer interaction[128] etc. A scene may be recorded as 3D point clouds using 3D scanning tools like LiDAR or RGB-D sensors. Unlike images, point clouds are sparse, disordered, and irregular, [120] makes scene interpretation extremely challenging.
In recent years, numerous deep learning methods for 3D scene understanding have been proposed, which can be divided into projection-based, voxel-based, and point-based methods. Concretely, projection-based methods (e.g., MV3D[91], PointPillars[92], MVCNN[93] ) project 3D points onto various image planes and employ 2D CNN-based backbones for feature extraction. Voxel-based methods convert point clouds into regular voxel grids to facilitate 3D convolution operations (e.g., VoxNet[94], SSCNet[95]), and some works improve their efficiency through sparse convolution (e.g., MinkowskiNet[96], SSCNs[97], Embodiedscan[98]). In contrast, point-based methods process point clouds directly (e.g., PointNet[99], PointNet++[100], PointMLP[101]). Recently, to achieve model scalability, Transformers-based (e.g., PointTransformer[102], Swin3d[103], PT2[104], PT3[105], 3D-VisTA[106], LEO[107], PQ3D[108]) and Mamba-based (e.g., PointMamba[109], PCM[110], Mamba3D[111]) architectures have emerged. It is worth noting that in addition to directly using features from point clouds, PQ3D[108] also seamlessly combines features from multi-view images and voxels to enhance scene understanding capabilities.
@@ IV-A3 Active Exploration
The previously introduced 3D scene understanding methods endow robots with the ability to perceive the environment in a passive manner. In such cases, the perception system’s information acquisition and decision-making do not adapt to the evolving scene. However, passive perception serves as a crucial foundation for active exploration. Given that robots are capable of movement and frequent interaction with their surroundings, they should also be able to explore and perceive their environment actively. The relationship between them is shown in Fig. 7. Current methods addressing active perception focus on interacting with the environment[112, 113] or by changing the viewing direction to obtain more visual information[114, 115, 116, 117].
@@
For example, Pinto et al.[112] proposed a curious robot that learns visual representations through physical interaction with the environment rather than relying solely on category labels in a dataset. To address the challenge of interactive object perception across robots with varying morphologies, Tatiya et al.[113] proposed a multi-stage projection framework that transfers implicit knowledge through learned exploratory interactions, enabling robots to effectively recognize object properties without the need to relearn from scratch. Recognizing the challenge of autonomously capturing informative observations, Jayaraman et al. [114] proposed a reinforcement learning method where an agent learns to actively acquire informative visual observations by reducing its uncertainty about unobserved parts of its environment, using recurrent neural networks for the active completion of panoramic scenes and 3D object shapes. NeU-NBV[115] introduced a mapless planning framework that iteratively positions an RGB camera to capture the most informative images of an unknown scene, using a novel uncertainty estimation in image-based neural rendering to guide data collection towards the most uncertain views. Hu et al.[116] developed a robot exploration algorithm that predicts the value of future states using a state value function, combining offline Monte-Carlo training, online Temporal Difference adaptation, and an intrinsic reward function based on sensor information coverage. To address the issue of accidental input in open-world environments, Fan et al.[117] treated active recognition as a sequential evidence-gathering process, providing step-by-step uncertainty quantification and reliable prediction under evidence combination theory while effectively characterizing the merit of actions in open-world environments through a specially developed reward function.
TABLE V: Comparison of Different 3D VG Methods.
Type Method Years Visual Input LLM-base
Two-stage ScanRefer[129] 2020 3D $\times$
ReferIt3D[130] 2020 3D $\times$
TGNN[131] 2021 3D $\times$
SAT[132] 2021 3D+2D $\times$
FFL-3DOG[133] 2021 3D $\times$
3DVG-Transformer[134] 2021 3D $\times$
LanguageRefer[135] 2022 3D $\times$
LAR[136] 2022 3D $\times$
MVT[137] 2022 3D $\times$
LLM-Grounder[138] 2023 3D $\surd$
ZSVG3D[139] 2023 3D $\surd$
GPS[140] 2024 3D $\times$
One-stage 3D-SPS[141] 2022 3D+2D $\times$
BUTD-DETR[142] 2022 3D $\times$
EDA[143] 2023 3D $\times$
ReGround3D[144] 2024 3D+2D $\surd$
IV-B 3D Visual Grounding
Unlike traditional 2D visual grounding (VG), which operates within the confines of flat images, 3D VG incorporates depth, perspective, and spatial relationships between objects, providing a more robust framework for agents to interact with their environment. The task of 3D VG involves locating objects within a 3D environment using natural language descriptions [129, 130]. As summarized in Table V, recent methodologies in 3D visual grounding can be roughly divided into two categories: two-stage and one-stage methods[145].
IV-B1 Two-stage 3D Visual Grounding methods
Similar to corresponding 2D tasks[146], early research in 3D grounding predominantly utilized a two-stage detect-then-match pipeline. They initially employ pretrained detector[147] or segmentor[148, 149, 150] to extract features from numerous object proposals within a 3D scene, which are then fused with linguistic query features to match the target object. The focus of the two-stage research is mainly on the second stage, such as exploring the correlation between object proposal features and linguistic query features to select the best-matched object. ReferIt3D [130] and TGNN [131] not only learn to match the proposal features with textual embedding but also encode the contextual relationship among the objects via graph neural networks. To enhance 3D visual grounding in free-form descriptions and irregular point cloud, FFL-3DOG [133] utilized a language scene graph for phrase correlations, a multi-level 3D proposal relation graph for enriching visual features, and a description-guided 3D visual graph for encoding global contexts.
Figure 8: The diagram of two-stage (upper) and one-stage (bottom) 3D visual grounding methods[141]. (a) shows the example of 3D visual grounding. (b) two-stage method includes Sparse proposals that may overlook the target in the detection stage and Dense proposals that may confuse the matching stage. (c) one-stage methods can progressively select keypoints (blue points $\longrightarrow$ red points $\longrightarrow$ green points) with the guidance of the language description.
Recently, as the transformer architecture has demonstrated outstanding performance in natural language processing [151, 152] and computer vision tasks [153, 14], research has increasingly focused on using transformers for extracting and fusing visual language features in 3D visual grounding tasks. For example, LanguageRefer [135] employed a transformer-based architecture combining 3D spatial embeddings, language descriptions, and class label embeddings to achieve robust 3D visual grounding. 3DVG-Transformer [134] is a relation-aware visual grounding method for 3D point clouds, featuring a coordinate-guided contextual aggregation module for relation-enhanced proposal generation and a multiplex attention module for cross-modal proposal disambiguation. To enable more fine-grained reasoning of 3D objects and referring expressions, TransRefer3D [154] enhanced cross-modal feature representation using entity-and-relation aware attention, incorporating self-attention, entity-aware attention and relation-aware attention. GPS[140] proposes a unified learning framework to distill knowledge from million-scale 3D vision-language dataset (i.e. SCENEVERSE[140]) by leveraging three levels of contrastive alignment learning and masked language modeling objective learning. Most of the above methods for 3D VG focus on specific viewpoints, but the learned visual-linguistic correspondences may fail when the viewpoint changes. In order to learn more view-robust visual representations, MVT [137] proposed a multi-view transformer that learns view-independent multi-modal representations.To mitigate the limitations of sparse, noisy, and incomplete point clouds, various methods have explored the incorporation of detailed 2D visual features from captured (e.g., SAT [132] or synthesized (e.g., LAR [136]) images to enhance 3D visual grounding tasks.

### v8

@@ III-B Real-Scene Based Simulators
Achieving universal embodied agents in household activities is a primary focus in embodied AI. These embodied agents need to deeply understand human daily life and perform complex embodied tasks such as navigation and interaction in indoor environments. To meet the demands of these complex tasks, the simulated environments need to be close to the real world, which places high demands on the complexity and realism of the simulators. These simulators mostly collect data from the real world, create photorealistic 3D assets, and build scenes using 3D game engines like UE5 and Unity. The rich and realistic scenes make simulators based on real world environments the top choice in household activities.
@@
SAPIEN [48] stands out for its design, specifically tailored to simulate interactions with joint objects like doors, cabinets, and drawers. VirtualHome [49] is notable for its unique environment graph, which facilitates high-level embodied planning based on natural language descriptions of environments. While AI2-ThOR [50] offers a wealth of interactive scenes, these interactions, similar to those in VirtualHome, are script-based and lack real physical interactions. This design suffices for embodied tasks not requiring fine-grained interactions. Both iGibson [51] and TDW [52] provide fine-grained embodied control and highly simulated physical interactions. iGibson excels in offering abundant and realistic large-scale scenes, making it suitable for complex and long-term mobile operations, whereas TDW allows greater user freedom in scene expansion and features unique audio and flexible fluid simulations, making it indispensable for related simulation scenarios. Matterport3D [53], a foundational 2D-3D visual dataset, is widely used and extended in embodied AI benchmarks. Although the embodied agent in Habitat lacks interaction capabilities, its extensive indoor scenes, user-friendly interfaces, and open framework make it highly regarded in embodied navigation. InfiniteWorld [54] focuses on unified and scalable simulation framework and implemented various improvements and the latest implicit asset reconstruction, as well as natural language-driven scene generation and editing. It provides strong support for complex robotic interactions through distributed collaboration, AI assistance, and Human-in-the-Loop.
@@
Besides, automated simulation scene construction is greatly beneficial for obtaining high-quality embodied data. RoboGen [55] customizes tasks from randomly sampled 3D assets through LLMs, thereby creating scenes and automatically training agents; HOLODECK [56] can automatically customize corresponding high-quality simulation scenes in AI2-THOR based on human instructions; PhyScene [57] generates interactive and physically consistent high-quality 3D scenes based on conditional diffusion. The Allen Institute for Artificial Intelligence expanded AI2-THOR and proposed ProcTHOR [58], which can automatically generate simulated scenes with sufficient interactivity, diversity, and rationality.
@@ IV-A Active Visual Perception
TABLE III: The comparison of the active visual perception methods.
@@
vSLAM Traditional vSLAM MonoSLAM[60], ORB-SLAM[61], LSD-SLAM[62]
Semantic vSLAM SLAM++[63], QuadricSLAM[64], So-SLAM[65],
SG-SLAM[66], OVD-SLAM[67], GS-SLAM[68]
@@
3D-VisTA[83], LEO[84], PQ3D[85], PointMamba[86], Mamba3D[87]
@@
Active visual perception systems require fundamental capabilities such as state estimation, scene perception, and environment exploration. As shown in Fig. 5, these capabilities have been extensively studied within the domains of Visual Simultaneous Localization and Mapping (vSLAM)[94, 95], 3D Scene Understanding[96], and Active Exploration[11]. These research areas contribute to developing robust active visual perception systems, facilitating improved environmental interaction and navigation in complex, dynamic settings. We briefly introduce these three components and summarize the methods mentioned in each part in Table III.
@@
Figure 5: The schematic diagram of active visual perception. Visual SLAM and 3D Scene Understanding provide the foundation for passive visual perception, while active exploration provides activeness to the passive perception system. These elements works collaboratively for the active visual perception system.
IV-A1 Visual Simultaneous Localization and Mapping
Simultaneous Localization and Mapping (SLAM) aims to determine a robot’s position within an unknown environment while simultaneously constructing a map of the environment [97]. Range-based SLAM [98, 99] relies on rangefinders, such as laser scanners, radar, and sonar, to generate point cloud representations. However, this approach is costly and provides limited environmental information. In contrast, Visual SLAM (vSLAM) [94, 95] employs on-board cameras to capture image frames and build environmental representations. Its advantages include low hardware costs, high accuracy in small-scale scenarios, and the ability to capture rich environmental details. Classical vSLAM can be broadly categorized into Traditional vSLAM and Semantic vSLAM [95]. Traditional vSLAM uses image data and multi-view geometry to estimate a robot’s pose and construct low-level maps (e.g., sparse, semi-dense, or dense point clouds) through methods like filter-based approaches (e.g., MonoSLAM [60]), keyframe-based methods (e.g., ORB-SLAM [61]), and direct tracking techniques (e.g., LSD-SLAM [62]). However, low-level maps do not directly correspond to objects, making them challenging for robots to interpret and utilize. Semantic vSLAM addresses this limitation by integrating semantic information, enhancing robots’ ability to perceive and navigate unexplored environments.
@@ IV-A1 Visual Simultaneous Localization and Mapping
3D scene understanding [100] aims to distinguish objects’ semantics, identify their locations, and infer the geometric attributes from 3D scene data [101], which is fundamental in autonomous driving[102], robot navigation[103], and human-computer interaction[104] etc. A scene may be recorded as 3D point clouds using 3D scanning tools like LiDAR or RGB-D sensors. Unlike images, point clouds are sparse, disordered, and irregular. Recent advances in deep learning for 3D scene understanding can be categorized into projection-based, voxel-based, and point-based methods. Concretely, projection-based methods (e.g., MV3D[69], PointPillars[70], MVCNN[71] ) project 3D points onto various image planes and employ 2D CNN-based backbones for feature extraction. Voxel-based methods convert point clouds into regular voxel grids to facilitate 3D convolution operations (e.g., VoxNet[72], SSCNet[73]), and some works improve their efficiency through sparse convolution (e.g., MinkowskiNet[74], SSCNs[75], Embodiedscan[76]). In contrast, point-based methods process point clouds directly (e.g., PointNet[77], PointNet++[78], PointMLP[79]). Recently, to achieve model scalability, Transformers-based (e.g., PointTransformer[80], Swin3d[81], PT2[82], 3D-VisTA[83], LEO[84], PQ3D[85] ) and Mamba-based (e.g., PointMamba[86], Mamba3D[87]) architectures have emerged. Notably, PQ3D[85] enhances scene understanding by seamlessly integrating features from point clouds, multi-view images, and voxels.
@@ IV-A3 Active Exploration
The 3D scene understanding methods allow robots to passively perceive the environment, with static information acquisition and decision-making regardless of scene changes. Thus, while passive perception is essential, it must be complemented by active exploration, enabling robots to dynamically interact with and perceive their surroundings. The relationship between them is shown in Fig. 5. Current methods addressing active perception focus on interacting with the environment[88, 89] or by changing the viewing direction to obtain more visual information[90, 91, 92, 93].
@@
For example, Pinto et al.[88] proposed a curious robot that learns visual representations through physical interaction with the environment rather than relying solely on dataset category labels. To address the challenge of interactive object perception across robots with varying morphologies, Tatiya et al.[89] proposed a multi-stage projection framework that transfers implicit knowledge through learned exploratory interactions. Recognizing the challenge of autonomously capturing informative observations, Jayaraman et al. [90] proposed a reinforcement learning method where an agent learns to actively acquire informative visual observations by reducing its uncertainty about unobserved parts of its environment. NeU-NBV[91] introduced a mapless planning framework that iteratively positions an RGB camera to capture the most informative images of an unknown scene. Hu et al.[92] developed a robot exploration algorithm that predicts the value of future states using a state value function. To address the issue of accidental input in open-world environments, Fan et al.[93] treated active recognition as a sequential evidence-gathering process, providing step-by-step uncertainty quantification and reliable prediction under evidence combination theory.
IV-B Visual Language Navigation